A Bibliography of Supercomputing ’2012

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

13 December 2012 Version 1.04

Abstract

This bibliography records articles presented at the Supercomputing ’2012 conference.

Title word cross-reference

1 [TPKP12]. 3 [NSM12]. N [INM12, MD12, SRK^{+}12, MCD^{+}12].

-body [INM12, MD12, SRK^{+}12]. -D [NSM12, TPKP12]. -point [MCD^{+}12].

100Gbps [HSH^{+}12]. ’12 [Hol12].

2.0 [NSM12]. 20.5Gflops [MD12]. 20.5Gflops/W [MD12]. 2012 [Hol12].

500km [HSH^{+}12].

8 [MD12].

Accelerating [CHA12]. acceleration [KVC^{+}12]. accelerator [MD12]. adaptive [BGI^{+}12, LYC^{+}12, PBA12]. adjoint [RMNM^{+}12]. advances [HPS^{+}12]. advection [NLL^{+}12]. aggregation [IMB^{+}12, KVC^{+}12]. algebra [GGDS^{+}12]. algebraic [SBB^{+}12]. algorithm [AML^{+}12, KGPH12, PPA^{+}12]. algorithms [DRL12]. alignment [KGPH12]. Alleviating [RFD^{+}12]. AMR [WLX^{+}12]. Analysis [Hol12, BAB^{+}12, BCR^{+}12, CS12, KVR12, LLG12, TW12, ZKW^{+}12]. analytics [JSL^{+}12, KN12]. analyze [BGI^{+}12]. aperture [PTS^{+}12]. Application [CKWC12, DFB^{+}12, ET12, LSG12, NA12, SZM^{+}12]. application-aware [ET12]. applications [BGL^{+}12, BGI^{+}12, CS12, KMHK12, LVY12a, ME12, NCB^{+}12, RLY^{+}12, SBG^{+}12]. applied [AML^{+}12]. approach [KN12, LSG12]. approximate [KGPH12]. architecture [CHA12, PWD^{+}12]. architectures [RMNM^{+}12]. archive [AMF^{+}12]. area [HSH^{+}12, RLY^{+}12]. arrhythmias [MRG^{+}12]. Aspen [SV12]. astronomical [MCD^{+}12]. astrophysical [INM12]. ATLAS [KVR12]. auto [JTD^{+}12]. auto-tuning [JTD^{+}12]. automated [LLG12]. Automatic [PCR12]. avoiding [GGDS^{+}12, LBDS12]. aware [ET12, FMR12, IMB^{+}12, SPK^{+}12, WHT^{+}12].

background [GSS12]. backprojection [PTS^{+}12]. backprojection-based [PTS^{+}12]. balancing [AML^{+}12]. Bamboo [NCB^{+}12]. barriers [CPW^{+}12]. based [CXB^{+}12, FBR^{+}12, IRJ^{+}12, LV12, PTS^{+}12, RBAC12, WLX^{+}12]. Bayesian [BTBG^{+}12, DKC12, NA12]. behavior [AMF^{+}12, LLG12]. benchmark [VSL12]. benefit [SBG^{+}12]. between [KMHK12]. beyond [LSG12]. BG [HMF^{+}12]. BG/Q [HMF^{+}12]. big [KN12]. Billion [CKS^{+}12]. Billion-particle [CKS^{+}12]. binary [LVY12a]. biomolecular [SZM^{+}12]. BitTorrent [DRL12]. blocking [SMM^{+}12]. Blue [CEH^{+}12, CKWC12, Doi12]. body [INM12, MD12, SRK^{+}12]. bottlenecks [CS12, XCD^{+}12]. breadth [BAP12]. breadth-first [BAP12]. Breaking [CPW^{+}12]. broadcasts [DRL12]. BT [VSL12]. Byte [JSL^{+}12]. Byte-precision [JSL^{+}12].

caches [AVG^{+}12]. can [SBG^{+}12]. Cascade [FBR^{+}12]. cell [WLX^{+}12]. cell-based [WLX^{+}12]. cellular [MRG^{+}12]. Center [Hol12]. centric [KN12]. Characterizing [OdSSP12, XCD^{+}12]. checkpointing [IMB^{+}12, RFD^{+}12, SMM^{+}12]. chromodynamics [Doi12]. circuits [LVY^{+}12b]. City [Hol12]. class [REP^{+}12].

Classifying [LVY12a]. closer [GCSK12]. cloud [DKC12, KN12, PBA12]. clouds [MJDN12]. cluster [CKS^{+}12]. clustering [DRL12]. clusters [RBAC12]. Code [REP^{+}12]. codes [JTD^{+}12]. coexistence [AVG^{+}12]. cognitive [PWD^{+}12]. coherence [AVG^{+}12]. collective [HS12]. collectives [BGL^{+}12]. Combining [BAB^{+}12]. Communication [GGDS^{+}12, LBDS12, SZM^{+}12, TPKP12]. Communication-avoiding [LBDS12]. communications [HS12]. communicators [BGL^{+}12]. Compass [PWD^{+}12]. Compiler [DZKS12, ME12]. Compiler-directed [DZKS12]. compressible [HRCK12]. compression [IMB^{+}12]. computation [MCD^{+}12, NLL^{+}12, PTS^{+}12, TW12]. computations [BPB12]. compute [DKC12]. computer [INM12]. Computing [Hol12, ET12, FME^{+}12, LV12, PWD^{+}12, VSL12]. Conference [Hol12]. conquer [MGK^{+}12]. constrained [MJDN12]. Containment [CLS^{+}12]. convenient [CSC12]. Convention [Hol12]. cooling [KN12]. core [AG12, IAF^{+}12, PTS^{+}12]. correction [FME^{+}12]. correlation [CKS^{+}12]. correlations [MCD^{+}12]. corruption [FME^{+}12]. cosmic [GSS12]. cosmology [WLX^{+}12]. Cost [MJDN12]. Cost- [MJDN12]. costs [KN12]. coupled [CHA12]. CPU [CHA12, RBAC12]. Cray [FBR^{+}12, SZM^{+}12]. Critical [CS12]. CUDA [LLG12]. cut [CXB^{+}12]. cut-through [CXB^{+}12].

D [NSM12, TPKP12]. Data [KBE^{+}12, AG12, BGI^{+}12, CKWC12, FME^{+}12, IMB^{+}12, KN12, KVC^{+}12, MCD^{+}12, MKB^{+}12, NCB^{+}12, PPA^{+}12, RLY^{+}12, SKCD12, TW12, ZKW^{+}12]. data-aware [IMB^{+}12]. data-centric [KN12]. data-driven [NCB^{+}12]. Data-intensive [KBE^{+}12, RLY^{+}12, SKCD12]. data-movement [AG12]. datacenter [KMHK12]. Dataflow [MMVK12]. Dataflow-driven [MMVK12]. datasets [KBE^{+}12]. DBSCAN [PPA^{+}12]. deadline [MJDN12]. deadline-constrained [MJDN12]. deadlock [HPS^{+}12]. deduplication [MKB^{+}12]. degradation [DFB^{+}12]. Demonstrating [HSH^{+}12]. demotions [WHT^{+}12]. Design [SMM^{+}12, SKYS12, SPK^{+}12, ZKW^{+}12, IRJ^{+}12, RLY^{+}12]. Designing [GKZ12]. detail [JSL^{+}12]. Detection [FME^{+}12, HPS^{+}12]. diagnosing [CS12]. direct [HRCK12]. directed [DZKS12]. Direction [BAP12]. Direction-optimizing [BAP12]. directive [LV12]. directive-based [LV12]. discovery [ET12]. disjoint [PPA^{+}12]. disjoint-set [PPA^{+}12]. distributed [CPW^{+}12, LMB12, REP^{+}12]. distributed-memory [CPW^{+}12]. divide [MGK^{+}12]. domain [SV12]. domains [CLS^{+}12]. Dragonfly [FBR^{+}12]. DRAM [SL12, WHT^{+}12]. driven [MMVK12, NCB^{+}12]. drug [MRG^{+}12]. drug-induced [MRG^{+}12]. DSPs [IAF^{+}12]. dynamic [WHT^{+}12].

Early [LV12]. earthquake [CSC12]. ECC [LYC^{+}12]. effective [AG12]. effectiveness [ET12]. efficiency [FMR12]. Efficient [DRL12, KVC^{+}12, PTS^{+}12, CLS^{+}12, LYC^{+}12, PBA12, SKCD12]. emerging [RMNM^{+}12, WKS^{+}12]. enable [BAB^{+}12, SPK^{+}12]. end [SKYS12]. end-to-end [SKYS12]. energy [FMR12, KN12, SKCD12]. energy-efficient [SKCD12]. ensembles [MJDN12]. equivalencing [LLG12]. error [HPS^{+}12, LVY12a]. estimating [DFB^{+}12]. evaluation [CSC12, LV12]. ever [ABR^{+}12]. exascale [CLS^{+}12, LV12, LSG12, VSL12]. execution [REP^{+}12]. exploration [CPW^{+}12]. expressing [BTSA12]. Extending [VSL12]. Extreme [BTBG^{+}12, BAB^{+}12, HMF^{+}12, LVY12a]. Extreme-scale [BTBG^{+}12, BAB^{+}12, LVY12a]. extremely [FSM^{+}12].

failures [SL12]. Fault [GCSK12]. FFT [NSM12, TPKP12]. field [SL12]. fields [NLL^{+}12]. file [DZKS12, LMB12]. filtering [KBE^{+}12]. fine [SZM^{+}12]. fine-grained [SZM^{+}12]. First [ABR^{+}12, BAP12]. First-ever [ABR^{+}12]. flexible [CLS^{+}12]. flow [NLL^{+}12]. flows [HRCK12, LLG12]. forests [SBB^{+}12]. form [NCB^{+}12]. Forward [RMNM^{+}12]. forwarding [CXB^{+}12]. fragment [AML^{+}12]. framework [JTD^{+}12, RBAC12, TPKP12]. friendly [CKS^{+}12]. FTLE [NLL^{+}12]. full [ABR^{+}12]. fused [TW12].

gene [NA12, CEH^{+}12, CKWC12, Doi12]. Gene/Q [CEH^{+}12, CKWC12, Doi12]. general [FSM^{+}12, IAF^{+}12]. general-purpose [IAF^{+}12]. generation [KVR12, PCR12, REP^{+}12]. geometric [SBB^{+}12, WKS^{+}12]. geometric-algebraic [SBB^{+}12]. Google [DKC12]. governed [BTBG^{+}12]. GPU [CHA12, LV12, MMVK12, NSM12, RBAC12, RMNM^{+}12]. GPUs [CSKH12]. grained [SZM^{+}12]. granularity [LYC^{+}12]. GRAPE [MD12]. GRAPE-8 [MD12]. graph [CPW^{+}12, FMR12, SKCD12]. gravitational [INM12, MD12]. grid [KVR12]. GridFTP [LVY^{+}12b].

Hardware [AVG^{+}12, SBG^{+}12].

Hardware-software [AVG^{+}12]. HDFS [IRJ^{+}12]. heart [MRG^{+}12]. heterogeneous [DRL12, GKZ12, PCR12]. Heuristic [AML^{+}12]. Hierarchical [WLX^{+}12, DZKS12]. High [BDE^{+}12, FSM^{+}12, HRCK12, Hol12, IRJ^{+}12, CSKH12, CSC12, FME^{+}12, IAF^{+}12]. High-performance [FSM^{+}12, CSKH12, CSC12, FME^{+}12]. hood [CEH^{+}12]. Host [DKC12]. HPC [CKS^{+}12, DFB^{+}12, FBR^{+}12, GCSK12, IAF^{+}12, MKB^{+}12]. human [MRG^{+}12]. Hybridizing [LSG12].

I/O [BCR^{+}12, KVC^{+}12]. IaaS [MJDN12]. IBM [CEH^{+}12, CKWC12]. implementation [LBDS12, SKYS12]. in-situ [BAB^{+}12]. in-transit [BAB^{+}12]. independence [BTSA12]. induced [MRG^{+}12]. InfiniBand [IRJ^{+}12, SPK^{+}12]. inflation [OdSSP12]. instrumentation [LVY12a]. intelligent [SKYS12]. intensive [KBE^{+}12, RLY^{+}12, SKCD12]. interest [MGK^{+}12]. interference [KMHK12]. International [Hol12]. interrupts [CXB^{+}12]. inverse [BTBG^{+}12]. irregular [REP^{+}12]. issues [RLY^{+}12, RFD^{+}12].

kernel [MMVK12].

Lake [Hol12]. language [SV12]. Large [SKCD12, AMF^{+}12, BGI^{+}12, CKS^{+}12, FME^{+}12, FSM^{+}12, KBE^{+}12, MCD^{+}12, RMNM^{+}12]. Large-scale [SKCD12, AMF^{+}12, BGI^{+}12, CKS^{+}12, FME^{+}12, FSM^{+}12, MCD^{+}12, RMNM^{+}12]. latency [NCB^{+}12]. latency-tolerant [NCB^{+}12]. lattice [Doi12]. layout [DZKS12]. learning [NA12]. Legion [BTSA12]. level [GSS12, JSL^{+}12, ME12]. linear [GGDS^{+}12]. live [KMHK12]. load [AML^{+}12, DKC12]. load-balancing [AML^{+}12]. local [AVG^{+}12]. locality [BTSA12]. lock [CS12]. logical [BTSA12]. look [GCSK12]. Looking [CEH^{+}12]. loops [REP^{+}12]. low [IAF^{+}12, TPKP12]. low-communication [TPKP12]. Lustre [HSH^{+}12].

machine [TW12]. machines [CPW^{+}12, GKZ12]. MAGE [LYC^{+}12]. making [GSS12]. management [ET12, WHT^{+}12, ZKW^{+}12]. Managing [AG12]. many [PTS^{+}12]. many-core [PTS^{+}12]. manycore [WKS^{+}12]. map [GSS12]. map-making [GSS12]. Mapping [BGL^{+}12, WLX^{+}12]. MapReduce [CHA12]. Massively [SLC^{+}12, SRK^{+}12]. matching [KGPH12]. maximize [BPB12]. McrEngine [IMB^{+}12]. Measuring [KMHK12]. memories [AVG^{+}12]. memory [AG12, CPW^{+}12, LYC^{+}12, REP^{+}12, SBG^{+}12]. method [AML^{+}12, DFB^{+}12]. microscope [GCSK12]. microwave [GSS12]. migrations [WHT^{+}12]. mining [FMR12]. mitigating [OdSSP12]. model [DKC12, GKZ12]. modeling [KVR12, MRG^{+}12, SMM^{+}12, SV12]. models [LV12]. molecular [AML^{+}12]. movement [AG12]. moving [LSG12]. MPI [HPS^{+}12, NCB^{+}12]. multi [HMF^{+}12, IAF^{+}12, JTD^{+}12, LSG12, MMVK12, NSM12, PBA12, WKS^{+}12]. multi- [WKS^{+}12]. multi-cloud [PBA12]. multi-core [IAF^{+}12]. multi-GPU [NSM12]. multi-kernel [MMVK12]. multi-objective [JTD^{+}12]. multi-petaflop [HMF^{+}12]. multi-petaflops [LSG12]. multicore [DFB^{+}12]. multigrid [SBB^{+}12, WKS^{+}12]. multiple [MGK^{+}12]. multithreaded [CS12, KGPH12]. MUST [HPS^{+}12].

NAS [VSL12]. NDGF [KVR12]. neighborhood [HS12]. network [CEH^{+}12, DRL12, FBR^{+}12, HSH^{+}12, KGPH12, NA12, SKYS12, SPK^{+}12]. network-topology-aware [SPK^{+}12]. Networking [Hol12, CXB^{+}12]. networks [BGL^{+}12, DRL12, NA12]. non [SMM^{+}12]. non-blocking [SMM^{+}12]. Novel [BGI^{+}12]. November [Hol12]. NUMA [FMR12]. NUMA-aware [FMR12]. numerical [GGDS^{+}12, HRCK12, KBE^{+}12]. numerically [CSKH12].

O [BCR^{+}12, KVC^{+}12]. objective [JTD^{+}12]. observable [ABR^{+}12]. octrees [SBB^{+}12]. OpenACC [LSG12]. optimistic [CXB^{+}12]. Optimization [HS12, WKS^{+}12, DZKS12]. Optimizing [CXB^{+}12, MCD^{+}12, SZM^{+}12, BAP12]. orbital [AML^{+}12]. out-of-core [AG12]. output [XCD^{+}12]. overlapping [GGDS^{+}12]. overlay [CXB^{+}12]. overlay-based [CXB^{+}12].

Parallel [BCR^{+}12, NA12, NLL^{+}12, SBB^{+}12, GSS12, JTD^{+}12, LMB12, LBDS12, OdSSP12, PPA^{+}12, PCR12, REP^{+}12, SLC^{+}12, SRK^{+}12, VSL12, ZKW^{+}12]. parallelism [BPB12]. parallelization [AG12]. parallelized [ME12]. Parametric [LLG12]. particle [BCR^{+}12, CKS^{+}12, NLL^{+}12]. path [SKCD12]. Patus [CSC12]. PDEs [BTBG^{+}12]. Performance [Hol12, BDE^{+}12, BGI^{+}12, CSKH12, CSC12, DFB^{+}12, FME^{+}12, FMR12, FSM^{+}12, IAF^{+}12, IRJ^{+}12, LBDS12, MD12, MMVK12, RLY^{+}12, SV12]. Peta [Doi12]. Peta-scale [Doi12]. petaflop [HMF^{+}12]. petaflops [LSG12]. Pflops [INM12]. phase [HRCK12]. PIDX [KVC^{+}12]. pipelines [PCR12]. placement [SPK^{+}12]. point [CKS^{+}12, MCD^{+}12]. Portable [ME12]. power [IAF^{+}12, LYC^{+}12, WHT^{+}12]. practical [DFB^{+}12]. precision [JSL^{+}12]. preconditioner [GSS12]. prediction [DKC12, GCSK12]. prefetching [CKWC12]. preparing [BDE^{+}12]. principles [HS12]. problem [INM12]. problems [BTBG^{+}12, FSM^{+}12]. Proceedings [Hol12]. processes [SPK^{+}12]. processing [BAB^{+}12, JSL^{+}12]. processors [DFB^{+}12, PTS^{+}12, WKS^{+}12]. productive [LV12]. programming [FSM^{+}12, GKZ12, LV12]. programs [LLG12, OdSSP12]. projection [MMVK12]. propagation [RMNM^{+}12]. protocol [AVG^{+}12]. Protocols [RLY^{+}12, RFD^{+}12]. provisioning [MJDN12]. purpose [IAF^{+}12].

Q [CEH^{+}12, CKWC12, Doi12, HMF^{+}12]. QoS [SKYS12]. quantum [Doi12].

races [LLG12]. radar [PTS^{+}12]. radiation [BDE^{+}12]. RAMZzz [WHT^{+}12]. rank [WHT^{+}12]. rank-aware [WHT^{+}12]. ray [SLC^{+}12]. RDMA [IRJ^{+}12]. RDMA-based [IRJ^{+}12]. real [MRG^{+}12]. real-time [MRG^{+}12]. reduction [KN12]. regions [BTSA12, MGK^{+}12]. reliable [DRL12]. resilience [CLS^{+}12]. resilient [LYC^{+}12]. resolution [MRG^{+}12]. resources [KVR12]. restructuring [KVC^{+}12]. runtime [HPS^{+}12].

S3D [LSG12]. Salt [Hol12]. scalability [CPW^{+}12, RFD^{+}12]. Scalable [NSM12, CSKH12, CLS^{+}12, FBR^{+}12, IMB^{+}12, PPA^{+}12, PWD^{+}12, SPK^{+}12, ZKW^{+}12]. scale [AMF^{+}12, BAB^{+}12, BGI^{+}12, BTBG^{+}12, CKS^{+}12, Doi12, FME^{+}12, FSM^{+}12, HMF^{+}12, LVY12a, MCD^{+}12, RMNM^{+}12, SKCD12]. Scalia [PBA12]. scaling [MGK^{+}12]. scattering [SLC^{+}12]. scheduling [RBAC12]. scheme [CLS^{+}12, PBA12]. scientific [AMF^{+}12, BAB^{+}12, ET12, LVY12a, MJDN12, SBG^{+}12]. scripting [ZKW^{+}12]. search [BAP12]. section [CS12, ME12]. section-level [ME12]. seismic [RMNM^{+}12]. self [ET12]. self-management [ET12]. semidefinite [FSM^{+}12]. service [SPK^{+}12]. set [PPA^{+}12]. SGI [TW12]. shared [AG12]. shared-memory [AG12]. silent [FME^{+}12]. SIMD [CKS^{+}12]. SIMD-friendly [CKS^{+}12]. simulation [ABR^{+}12, BCR^{+}12, HMF^{+}12, INM12, KBE^{+}12, MD12, MRG^{+}12, SZM^{+}12]. simulations [BDE^{+}12, CSC12, HRCK12, MGK^{+}12, RMNM^{+}12, SLC^{+}12, WLX^{+}12]. simulator [PWD^{+}12]. situ [BAB^{+}12]. sky [HMF^{+}12]. soft [LVY12a]. software [AVG^{+}12, HRCK12, PCR12]. solver [CSKH12, FSM^{+}12, SRK^{+}12]. solvers [AG12]. space [SRK^{+}12]. space-time [SRK^{+}12]. sparse [AG12]. spatial [KBE^{+}12]. specific [SV12]. speed [CPW^{+}12]. stable [CSKH12]. static [AML^{+}12]. stencil [BPB12]. stencils [CSC12]. Storage [Hol12, DZKS12, MKB^{+}12, PBA12]. Strassen [LBDS12]. strategy [MGK^{+}12]. structure [NA12, PPA^{+}12]. study [MKB^{+}12, SL12]. sub [BGL^{+}12]. sub-communicators [BGL^{+}12]. supercomputer [CKWC12, Doi12, NSM12, XCD^{+}12]. supercomputing [SKCD12]. symbolic [LLG12]. synthetic [PTS^{+}12]. system [FBR^{+}12, IMB^{+}12, SMM^{+}12, SKYS12]. systems [CKS^{+}12, CLS^{+}12, DZKS12, ET12, GCSK12, LMB12, LYC^{+}12, MKB^{+}12, PCR12, REP^{+}12].

T* [KN12]. task [OdSSP12, WLX^{+}12]. techniques [FMR12]. throughput [HRCK12]. Tiling [BPB12]. time [MRG^{+}12, NLL^{+}12, OdSSP12, SRK^{+}12]. time-varying [NLL^{+}12]. Titan [BDE^{+}12]. tolerant [NCB^{+}12]. tomography [DRL12]. tool [LVY12a]. topology [SPK^{+}12]. torus [BGL^{+}12]. transactional [SBG^{+}12]. transfers [LVY^{+}12b]. transformations [MMVK12]. transit [BAB^{+}12]. translating [NCB^{+}12]. transport [BDE^{+}12]. traversal [SKCD12]. tree [LMB12]. tridiagonal [CSKH12]. trillion [BCR^{+}12, INM12]. trillion-body [INM12]. TSUBAME [NSM12]. tuning [JTD^{+}12, ME12]. two [CKS^{+}12, GSS12, HRCK12]. two-level [GSS12]. two-phase [HRCK12]. two-point [CKS^{+}12].

unified [GKZ12]. universe [ABR^{+}12, HMF^{+}12]. Unleashing [IAF^{+}12]. unstructured [SBB^{+}12]. UQ [BTBG^{+}12]. USA [Hol12]. Usage [AMF^{+}12]. using [CSKH12, DRL12, IMB^{+}12, LSG12, LVY12a, LVY^{+}12b, PPA^{+}12]. UT [Hol12]. UV2 [TW12].

value [RBAC12]. value-based [RBAC12]. ValuePack [RBAC12]. variable [JSL^{+}12]. varying [NLL^{+}12]. ventricles [MRG^{+}12]. via [KGPH12]. views [BGI^{+}12]. virtual [CXB^{+}12, LVY^{+}12b]. visualization [BCR^{+}12]. volunteer [ET12]. vulnerabilities [LVY12a].

W [MD12]. walk [LMB12]. wave

[RMNM^{+}12]. weather [MGK^{+}12]. wide [HSH^{+}12, RLY^{+}12]. wide-area [RLY^{+}12]. work [OdSSP12]. workflow [MJDN12]. workload [KVR12]. workloads [DFB^{+}12].

X [SLC^{+}12]. X-ray [SLC^{+}12]. XK6 [SZM^{+}12].

References

Alimi:2012:FEF

[ABR^{+}12] Jean-Michel Alimi, Vincent Bouillot, Yann Rasera, Vincent Reverdy, Pier-Stefano Corasaniti, Irène Balmès, Stéphane Requena, Xavier Delaruelle, and Jean-Noel Richet. First-ever full observable universe simulation. In Hollingsworth [Hol12], pages 73:1–73:11. ISBN 1-4673-0804-8. URL http:/

                                                    /conferences.computer.org/
                                                    sc/2012/papers/1000a034.pdf.

Avron:2012:MDM

[AG12] Haim Avron and Anshul Gupta. Managing data-movement for effective shared-memory parallelization of out-of-core sparse solvers. In Hollingsworth [Hol12], pages 102:1–102:11. ISBN 1-4673-0804-8. URL

http://conferences.computer.
org/sc/2012/papers/1000a062.
pdf.

Adams:2012:UBL

[AMF^{+}12] Ian F. Adams, Brian A. Madden, Joel C. Frank, Mark W. Storer, Ethan L. Miller, and Gene Harano. Usage behavior of a large-scale scientific archive. In Hollingsworth [Hol12], pages 86:1–86:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a017.pdf.

Alexeev:2012:HSL

[AML^{+}12] Yuri Alexeev, Ashutosh Mahajan, Sven Leyffer, Graham Fletcher, and Dmitri G. Fedorov. Heuristic static loadbalancing algorithm applied to the fragment molecular orbital method. In Hollingsworth [Hol12], pages 56:1–56:13. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a087.pdf.

Alvarez:2012:HSC

[AVG^{+}12] Lluc Alvarez, Lluís Vilanova, Marc Gonzalez, Xavier Martorell, Nacho Navarro, and Eduard Ayguade. Hardwaresoftware coherence protocol for the coexistence of caches and local memories. In Hollingsworth [Hol12], pages 89:1–89:11. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a074.pdf.

Bennett:2012:CST

[BAB^{+}12] Janine C. Bennett, Hasan Abbasi, Peer-Timo Bremer, Ray Grout, Attila Gyulassy, Tong Jin, Scott Klasky, Hemanth Kolla, Manish Parashar, Valerio Pascucci, Philippe Pebay, David Thompson, Hongfeng Yu, Fan Zhang, and Jacqueline Chen. Combining insitu and in-transit processing to enable extreme-scale scientific analysis. In Hollingsworth [Hol12], pages 49:1–49:9. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a089.pdf.

Beamer:2012:DOB

[BAP12] Scott Beamer, Krste Asanović, and David Patterson. Directionoptimizing breadth-first search. In Hollingsworth [Hol12], pages 12:1–12:10. ISBN 1-4673-0804-

                                        8. URL http://conferences.
                                        computer.org/sc/2012/papers/
                                        1000a019.pdf.

Byna:2012:PAV

[BCR^{+}12] Surendra Byna, Jerry Chou, Oliver Rübel, Prabhat, Homa Karimabadi, William S. Daughton, Vadim Roytershteyn, E. Wes Bethel, Mark Howison, Ke-Jou Hsu, Kuan-Wu Lin, Arie Shoshani, Andrew Uselton, and Kesheng Wu. Parallel I/O, analysis, and visualization of a trillion particle simulation. In Hollingsworth [Hol12], pages 59:1–59:12. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a103.pdf.

Baker:2012:HPR

[BDE^{+}12] C. Baker, G. Davidson, T. M.

Evans, S. Hamilton, J. Jarrell, and W. Joubert. High performance radiation transport simulations: preparing for Titan. In Hollingsworth [Hol12], pages 47:1–47:10. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a069.pdf.

Bhatele:2012:NVP

[BGI^{+}12] Abhinav Bhatele, Todd Gamblin, Katherine E. Isaacs, Brian T. N. Gunney, Martin Schulz, Peer-Timo Bremer, and Bernd Hamann. Novel views of performance data to analyze large-scale adaptive applications. In Hollingsworth [Hol12], pages 31:1–31:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a011.pdf.

Bhatele:2012:MAC

[BGL^{+}12] Abhinav Bhatele, Todd Gamblin, Steven H. Langer, Peer-Timo Bremer, Erik W. Draeger, Bernd Hamann, Katherine E. Isaacs, Aaditya G. Landge, Joshua A. Levine, Valerio Pascucci, Martin Schulz, and Charles H. Still. Mapping applications with collectives over sub-communicators on torus networks. In Hollingsworth [Hol12], pages 97:1–97:11. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a027.pdf.

Bandishti:2012:TSC

[BPB12] Vinayaka Bandishti, Irshad Pananilath, and Uday Bondhugula. Tiling stencil computations to maximize parallelism. In Hollingsworth [Hol12], pages 40:1–40:11. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a031.pdf.

Bui-Thanh:2012:ESU

[BTBG^{+}12] Tan Bui-Thanh, Carsten Burstedde, Omar Ghattas, James Martin, Georg Stadler, and Lucas C. Wilcox. Extremescale UQ for Bayesian inverse problems governed by PDEs. In Hollingsworth [Hol12], pages 3:1–3:11. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a003.pdf.

Bauer:2012:LEL

[BTSA12] Michael Bauer, Sean Treichler, Elliott Slaughter, and Alex Aiken. Legion: expressing locality and independence with logical regions. In Hollingsworth [Hol12], pages 66:1–66:11. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a065.pdf.

Chen:2012:LUH

[CEH^{+}12] Dong Chen, Noel Eisley, Philip Heidelberger, Sameer Kumar, Amith Mamidala, Fabrizio Petrini, Robert Senger, Yutaka Sugawara, Robert Walkup, Burkhard Steinmacher-Burow, Anamitra Choudhury, Yogish Sabharwal, Swati Singhal, and Jeffrey J. Parker. Looking under the hood of the IBM Blue Gene/Q network. In Hollingsworth [Hol12], pages 69:1–69:12. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a078.pdf.

Chen:2012:AMC

[CHA12] Linchuan Chen, Xin Huo, and Gagan Agrawal. Accelerating MapReduce on a coupled CPU–GPU architecture. In Hollingsworth [Hol12], pages 25:1–25:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a050.pdf.

Chhugani:2012:BPS

[CKS^{+}12] Jatin Chhugani, Changkyu Kim, Hemant Shukla, Jongsoo Park, Pradeep Dubey, John Shalf, and Horst D. Simon. Billion-particle SIMDfriendly two-point correlation on large-scale HPC cluster systems. In Hollingsworth [Hol12], pages 1:1–1:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a002.pdf.

Chung:2012:ADP

[CKWC12] I-Hsin Chung, Changhoan Kim, Hui-Fang Wen, and Guojing Cong. Application data prefetching on the IBM Blue Gene/Q supercomputer. In Hollingsworth [Hol12], pages 88:1–88:8. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a075.pdf.

Chung:2012:CDS

[CLS^{+}12] Jinsuk Chung, Ikhwan Lee, Michael Sullivan, Jee Ho Ryoo, Dong Wan Kim, Doe Hyun Yoon, Larry Kaplan, and Mattan Erez. Containment domains: a scalable, efficient, and flexible resilience scheme for exascale systems. In Hollingsworth [Hol12], pages 58:1–58:11. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a098.pdf.

Checconi:2012:BSS

[CPW^{+}12] Fabio Checconi, Fabrizio Petrini, Jeremiah Willcock, Andrew Lumsdaine, Anamitra Roy Choudhury, and Yogish Sabharwal. Breaking the speed and scalability barriers for graph exploration on distributedmemory machines. In Hollingsworth [Hol12], pages 13:1–13:12. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a020.pdf.

Chen:2012:CLA

[CS12] Guancheng Chen and Per Stenstrom. Critical lock analysis: diagnosing critical section bottlenecks in multithreaded applications. In Hollingsworth [Hol12], pages 71:1–71:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a099.pdf.

Christen:2012:PCH

[CSC12] Matthias Christen, Olaf Schenk, and Yifeng Cui. Patus for convenient high-performance stencils: evaluation in earthquake simulations. In Hollingsworth [Hol12], pages 11:1–11:10. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a014.pdf.

Chang:2012:SNS

[CSKH12] Li-Wen Chang, John A. Stratton, Hee-Seok Kim, and Wen-Mei W. Hwu. A scalable, numerically stable, highperformance tridiagonal solver using GPUs. In Hollingsworth [Hol12], pages 27:1–27:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a071.pdf.

Cui:2012:OOB

[CXB^{+}12] Zheng Cui, Lei Xia, Patrick G.

Bridges, Peter A. Dinda, and John R. Lange. Optimizing overlay-based virtual networking through optimistic interrupts and cut-through forwarding. In Hollingsworth [Hol12], pages 99:1–99:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a029.pdf.

Dwyer:2012:PME

[DFB^{+}12] Tyler Dwyer, Alexandra Fedorova, Sergey Blagodurov, Mark Roth, Fabien Gaud, and Jian Pei. A practical method for estimating performance degradation on multicore processors, and its application to HPC workloads. In Hollingsworth [Hol12], pages 83:1–83:11. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a093.pdf.

Di:2012:HLP

[DKC12] Sheng Di, Derrick Kondo, and Walfredo Cirne. Host load prediction in a Google compute cloud with a Bayesian model. In Hollingsworth [Hol12], pages 21:1–21:11. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a025.pdf.

Doi:2012:PSL

[Doi12]  Jun Doi.  Peta-scale lat-

tice quantum chromodynamics on a Blue Gene/Q supercomputer. In Hollingsworth [Hol12], pages 45:1–45:10. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a068.pdf.

Dichev:2012:ERN

[DRL12] Kiril Dichev, Fergal Reid, and Alexey Lastovetsky. Efficient and reliable network tomography in heterogeneous networks using BitTorrent broadcasts and clustering algorithms. In Hollingsworth [Hol12], pages 36:1–36:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a060.pdf.

Ding:2012:CDF

[DZKS12] Wei Ding, Yuanrui Zhang, Mahmut Kandemir, and Seung Woo Son. Compilerdirected file layout optimization for hierarchical storage systems. In Hollingsworth [Hol12], pages 41:1–41:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a030.pdf.

Estrada:2012:EAA

[ET12] Trilce Estrada and Michela Taufer. On the effectiveness of application-aware selfmanagement for scientific discovery in volunteer computing systems. In Hollingsworth [Hol12], pages 80:1–80:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a057.pdf.

Faanes:2012:CCS

[FBR^{+}12] Greg Faanes, Abdulla Bataineh, Duncan Roweth, Tom Court, Edwin Froese, Bob Alverson, Tim Johnson, Joe Kopnick, Mike Higgins, and James Reinhard. Cray Cascade: a scalable HPC system based on a Dragonfly network. In Hollingsworth [Hol12], pages 103:1–103:9. ISBN 1-4673-0804-

8. URL http://conferences.
computer.org/sc/2012/papers/
1000a079.pdf.

Fiala:2012:DCS

[FME^{+}12] David Fiala, Frank Mueller, Christian Engelmann, Rolf Riesen, Kurt Ferreira, and Ron Brightwell. Detection and correction of silent data corruption for large-scale high-performance computing. In Hollingsworth [Hol12], pages 78:1–78:12. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a046.pdf.

Frasca:2012:NAG

[FMR12] Michael Frasca, Kamesh Madduri, and Padma Raghavan. NUMA-aware graph mining techniques for performance and energy efficiency. In Hollingsworth [Hol12], pages 95:1–95:11. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a096.pdf.

Fujisawa:2012:HPG

[FSM^{+}12] Katsuki Fujisawa, Hitoshi Sato, Satoshi Matsuoka, Toshio Endo, Makoto Yamashita, and Maho Nakata. High-performance general solver for extremely largescale semidefinite programming problems. In Hollingsworth [Hol12], pages 93:1–93:11. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a082.pdf.

Gainaru:2012:FPU

[GCSK12] Ana Gainaru, Franck Cappello, Marc Snir, and William Kramer. Fault prediction under the microscope: a closer look into HPC systems. In Hollingsworth [Hol12], pages 77:1–77:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a048.pdf.

Georganas:2012:CAO

[GGDS^{+}12] Evangelos Georganas, Jorge González-Domínguez, Edgar Solomonik, Yili Zheng, Juan Touriño, and Katherine Yelick. Communication avoiding and overlapping for numerical linear algebra. In Hollingsworth [Hol12], pages 100:1–100:11. ISBN 1-4673-0804-8. URL

http://conferences.computer.
org/sc/2012/papers/1000a061.
pdf.

Garland:2012:DUP

[GKZ12] Michael Garland, Manjunath Kudlur, and Yili Zheng. Designing a unified programming model for heterogeneous machines. In Hollingsworth [Hol12], pages 67:1–67:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a064.pdf.

Grigori:2012:PTL

[GSS12] Laura Grigori, Radek Stompor, and Mikolaj Szydlarski. A parallel two-level preconditioner for cosmic microwave background map-making. In Hollingsworth [Hol12], pages 91:1–91:10. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a084.pdf.

Habib:2012:UES

[HMF^{+}12] Salman Habib, Vitali Morozov, Hal Finkel, Adrian Pope, Katrin Heitmann, Kalyan Kumaran, Tom Peterka, Joe Insley, David Daniel, Patricia Fasel, Nicholas Frontiere, and Zarija Lukić. The universe at extreme scale: multi-petaflop sky simulation on the BG/Q. In Hollingsworth [Hol12], pages 4:1–4:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a004.pdf.

Hollingsworth:2012:SPI

[Hol12] Jeffrey Hollingsworth, editor. SC ’12: Proceedings of the International Conference on High Performance Computing, Networking, Storage and Analysis, Salt Lake Convention Center, Salt Lake City, UT, USA, November 10–16, 2012. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2012. ISBN 1-4673-0804-8.

Hilbrich:2012:MRE

[HPS^{+}12] Tobias Hilbrich, Joachim Protze, Martin Schulz, Bronis R. de Supinski, and Matthias S. Müller. MPI runtime error detection with MUST: advances in deadlock detection. In Hollingsworth [Hol12], pages 30:1–30:10. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a010.pdf.

Hejazialhosseini:2012:HTS

[HRCK12] Babak Hejazialhosseini, Diego Rossinelli, Christian Conti, and Petros Koumoutsakos. High throughput software for direct numerical simulations of compressible two-phase flows. In Hollingsworth [Hol12], pages 16:1–16:12. ISBN 1-4673-0804-

8. URL http://conferences.
computer.org/sc/2012/papers/
1000a039.pdf.

Hoefler:2012:OPC

[HS12] Torsten Hoefler and Timo Schneider. Optimization principles for collective neighborhood communications. In Hollingsworth [Hol12], pages 98:1–98:10. ISBN 1-4673-0804-

8. URL http://conferences.
computer.org/sc/2012/papers/
1000a028.pdf.

Henschel:2012:DLW

[HSH^{+}12] Robert Henschel, Stephen Simms, David Hancock, Scott Michael, Tom Johnson, Nathan Heald, Thomas William, Donald Berry, Matt Allen, Richard Knepper, Matthew Davy, Matthew Link, and Craig A. Stewart. Demonstrating Lustre over a 100Gbps wide area network of 3,500km. In Hollingsworth [Hol12], pages 6:1–6:8. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a006.pdf.

Igual:2012:UHP

[IAF^{+}12] Francisco D. Igual, Murtaza Ali, Arnon Friedmann, Eric Stotzer, Timothy Wentz, and Robert A. van de Geijn. Unleashing the high performance and low power of multi-core DSPs for generalpurpose HPC. In Hollingsworth [Hol12], pages 26:1–26:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a070.pdf.

Islam:2012:MSC

[IMB^{+}12] Tanzima Zerin Islam, Kathryn Mohror, Saurabh Bagchi, Adam Moody, Bronis R. de Supinski, and Rudolf Eigenmann. McrEngine: a scalable checkpointing system using dataaware aggregation and compression. In Hollingsworth [Hol12], pages 17:1–17:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a023.pdf.

Ishiyama:2012:PAB

[INM12] Tomoaki Ishiyama, Keigo Nitadori, and Junichiro Makino. 4.45 pflops astrophysical Nbody simulation on K computer: the gravitational trillionbody problem. In Hollingsworth [Hol12], pages 5:1–5:10. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a005.pdf.

Islam:2012:HPR

[IRJ^{+}12] N. S. Islam, M. W. Rahman, J. Jose, R. Rajachandrasekar, H. Wang, H. Subramoni, C. Murthy, and D. K. Panda. High performance RDMA-based design of HDFS over Infini-Band. In Hollingsworth [Hol12], pages 35:1–35:12. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a058.pdf.

Jenkins:2012:BPL

[JSL^{+}12] John Jenkins, Eric R. Schendel, Sriram Lakshminarasimhan, David A. Boyuka II, Terry Rogers, Stephane Ethier, Robert Ross, Scott Klasky, and Nagiza F. Samatova. Byteprecision level of detail processing for variable precision analytics. In Hollingsworth [Hol12], pages 48:1–48:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a088.pdf.

Jordan:2012:MOA

[JTD^{+}12] Herbert Jordan, Peter Thoman, Juan J. Durillo, Simone Pellegrini, Philipp Gschwandtner, Thomas Fahringer, and Hans Moritsch. A multi-objective auto-tuning framework for parallel codes. In Hollingsworth [Hol12], pages 10:1–10:12. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a013.pdf.

Kanov:2012:DIS

[KBE^{+}12] Kalin Kanov, Randal Burns, Greg Eyink, Charles Meneveau, and Alexander Szalay. Data-intensive spatial filtering in large numerical simulation datasets. In Hollingsworth [Hol12], pages 60:1–60:9. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a101.pdf.

Khan:2012:MAN

[KGPH12] Arif M. Khan, David F. Gleich, Alex Pothen, and Mahantesh Halappanavar. A multithreaded algorithm for network alignment via approximate matching. In Hollingsworth [Hol12], pages 64:1–64:11. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a054.pdf.

Kambadur:2012:MIB

[KMHK12] Melanie Kambadur, Tipp Moseley, Rick Hank, and Martha A. Kim. Measuring interference between live datacenter applications. In Hollingsworth [Hol12], pages 51:1–51:12. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a036.pdf.

Kaushik:2012:DCC

[KN12] Rini T. Kaushik and Klara Nahrstedt. T*: a data-centric cooling energy costs reduction approach for big data analytics cloud. In Hollingsworth [Hol12], pages 52:1–52:11. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a037.pdf.

Kumar:2012:EDR

[KVC^{+}12] Sidharth Kumar, Venkatram Vishwanath, Philip Carns, Joshua A. Levine, Robert Latham, Giorgio Scorzelli, Hemanth Kolla, Ray Grout, Robert Ross, Michael E. Papka, Jacqueline Chen, and Valerio Pascucci. Efficient data restructuring and aggregation for I/O acceleration in PIDX. In Hollingsworth [Hol12], pages 50:1–50:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a090.pdf.

Karpenko:2012:AGW

[KVR12] Dmytro Karpenko, Roman Vitenberg, and Alexander L. Read. ATLAS grid workload on NDGF resources: analysis, modeling, and workload generation. In Hollingsworth [Hol12], pages 79:1–79:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a055.pdf.

Lipshitz:2012:CAP

[LBDS12] Benjamin Lipshitz, Grey Ballard, James Demmel, and Oded Schwartz. Communicationavoiding parallel Strassen: implementation and performance. In Hollingsworth [Hol12], pages 101:1–101:11. ISBN 1-4673-

0804-8. URL http://conferences.
computer.org/sc/2012/papers/
1000a063.pdf.

Li:2012:PFA

[LLG12] Peng Li, Guodong Li, and Ganesh Gopalakrishnan. Parametric flows: automated behavior equivalencing for symbolic analysis of races in CUDA programs. In Hollingsworth [Hol12], pages 29:1–29:10. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a009.pdf.

LaFon:2012:DFT

[LMB12] Jharrod LaFon, Satyajayant Misra, and Jon Bringhurst. On distributed file tree walk of parallel file systems. In Hollingsworth [Hol12], pages 87:1–87:11. ISBN 1-4673-0804-

                                        8. URL http://conferences.
                                        computer.org/sc/2012/papers/
                                        1000a015.pdf.

Levesque:2012:HEA

[LSG12] John M. Levesque, Ramanan Sankaran, and Ray Grout. Hybridizing S3D into an exascale application using OpenACC: an approach for moving to multi-petaflops and beyond. In Hollingsworth [Hol12], pages 15:1–15:11. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a040.pdf.

Lee:2012:EED

[LV12] Seyong Lee and Jeffrey S. Vetter. Early evaluation of directive-based GPU programming models for productive exascale computing. In Hollingsworth [Hol12], pages 23:1–23:11. ISBN 1-4673-0804-

                                        8. URL http://conferences.
                                        computer.org/sc/2012/papers/
                                        1000a051.pdf.

Li:2012:CSE

[LVY12a] Dong Li, Jeffrey S. Vetter, and Weikuan Yu. Classifying soft error vulnerabilities in extreme-scale scientific applications using a binary instrumentation tool. In Hollingsworth [Hol12], pages 57:1–57:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a097.pdf.

Liu:2012:UVC

[LVY^{+}12b] Z. Liu, M. Veeraraghavan, Z. Yan, C. Tracy, J. Tie, I. Foster, J. Dennis, J. Hick, Y. Li, and W. Yang. On using virtual circuits for GridFTP transfers. In Hollingsworth [Hol12], pages 81:1–81:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a056.pdf.

Li:2012:MAG

[LYC^{+}12] Sheng Li, Doe Hyun Yoon, Ke Chen, Jishen Zhao, Jung Ho Ahn, Jay B. Brockman, Yuan Xie, and Norman P. Jouppi. MAGE: adaptive granularity and ECC for resilient and power efficient memory systems. In Hollingsworth [Hol12], pages 33:1–33:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a041.pdf.

March:2012:OCP

[MCD^{+}12] William B. March, Kenneth Czechowski, Marat Dukhan, Thomas Benson, Dongryeol Lee, Andrew J. Connolly, Richard Vuduc, Edmond Chow, and Alexander G. Gray. Optimizing the computation of n-point correlations on large-scale astronomical data. In Hollingsworth [Hol12], pages 74:1–74:12. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a033.pdf.

Makino:2012:GAG

[MD12] Junichiro Makino and Hiroshi Daisaka. GRAPE-8: an accelerator for gravitational N-body simulation with 20.5Gflops/W performance. In Hollingsworth [Hol12], pages 104:1–104:10. ISBN 1-4673-0804-8. URL

                                        http://conferences.computer.
                                        org/sc/2012/papers/1000a081.
                                        pdf.

Mustafa:2012:PSL

[ME12] Dheya Mustafa and Rudolf Eigenmann. Portable sectionlevel tuning of compiler parallelized applications. In Hollingsworth [Hol12], pages 9:1–9:11. ISBN 1-4673-0804-

                                        8. URL http://conferences.
                                        computer.org/sc/2012/papers/
                                        1000a012.pdf.

Malakar:2012:DCS

[MGK^{+}12] Preeti Malakar, Thomas George, Sameer Kumar, Rashmi Mittal, Vijay Natarajan, Yogish Sabharwal, Vaibhav Saxena, and Sathish S. Vadhiyar. A divide and conquer strategy for scaling weather simulations with multiple regions of interest. In Hollingsworth [Hol12], pages 37:1–37:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a105.pdf.

Malawski:2012:CDC

[MJDN12] Maciej Malawski, Gideon Juve, Ewa Deelman, and Jarek Nabrzyski. Cost- and deadlineconstrained provisioning for scientific workflow ensembles in IaaS clouds. In Hollingsworth [Hol12], pages 22:1–22:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a024.pdf.

Meister:2012:SDD

[MKB^{+}12] Dirk Meister, Jürgen Kaiser, Andre Brinkmann, Toni Cortes, Michael Kuhn, and Julian Kunkel. A study on data deduplication in HPC storage systems. In Hollingsworth [Hol12], pages 7:1–7:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a008.pdf.

Meng:2012:DDG

[MMVK12] Jiayuan Meng, Vitali A. Morozov, Venkatram Vishwanath, and Kalyan Kumaran. Dataflowdriven GPU performance projection for multi-kernel transformations. In Hollingsworth [Hol12], pages 82:1–82:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a092.pdf.

Mirin:2012:TRT

[MRG^{+}12] Arthur A. Mirin, David F.

Richards, James N. Glosli, Erik W. Draeger, Bor Chan, Jean luc Fattebert, William D. Krauss, Tomas Oppelstrup, John Jeremy Rice, John A. Gunnels, Viatcheslav Gurev, Changhoan Kim, John Magerlein, Matthias Reumann, and Hui-Fang Wen. Toward realtime modeling of human heart ventricles at cellular resolution: simulation of drug-induced arrhythmias. In Hollingsworth [Hol12], pages 2:1–2:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a001.pdf.

Nikolova:2012:PBN

[NA12] Olga Nikolova and Srinivas Aluru. Parallel Bayesian network structure learning with application to gene networks. In Hollingsworth [Hol12], pages 63:1–63:9. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a052.pdf.

Nguyen:2012:BTM

[NCB^{+}12] Tan Nguyen, Pietro Cicotti, Eric Bylaska, Dan Quinlan, and Scott B. Baden. Bamboo: translating MPI applications to a latency-tolerant, datadriven form. In Hollingsworth [Hol12], pages 39:1–39:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a032.pdf.

Nouanesengsy:2012:PPA

[NLL^{+}12] Boonthanome Nouanesengsy, Teng-Yok Lee, Kewei Lu, Han-Wei Shen, and Tom Peterka. Parallel particle advection and FTLE computation for time-varying flow fields. In Hollingsworth [Hol12], pages 61:1–61:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a102.pdf.

Nukada:2012:SMG

[NSM12] Akira Nukada, Kento Sato, and Satoshi Matsuoka. Scalable multi-GPU 3-D FFT for TSUBAME 2.0 supercomputer. In Hollingsworth [Hol12], pages 44:1–44:10. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a044.pdf.

Olivier:2012:CMW

[OdSSP12] Stephen L. Olivier, Bronis R.

de Supinski, Martin Schulz, and Jan F. Prins. Characterizing and mitigating work time inflation in task parallel programs. In Hollingsworth [Hol12], pages 65:1–65:12. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a066.pdf.

Papaioannou:2012:SAS

[PBA12] Thanasis G. Papaioannou, Nicolas Bonvin, and Karl Aberer. Scalia: an adaptive scheme for efficient multi-cloud storage. In Hollingsworth [Hol12], pages 20:1–20:10. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a026.pdf.

Pienaar:2012:AGS

[PCR12] Jacques A. Pienaar, Srimat Chakradhar, and Anand Raghunathan. Automatic generation of software pipelines for heterogeneous parallel systems. In Hollingsworth [Hol12], pages 24:1–24:12. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a049.pdf.

Patwary:2012:NSP

[PPA^{+}12] Mostofa Ali Patwary, Diana Palsetia, Ankit Agrawal, Wei keng Liao, Fredrik Manne, and Alok Choudhary. A new scalable parallel DBSCAN algorithm using the disjoint-set data structure. In Hollingsworth [Hol12], pages 62:1–62:11. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a053.pdf.

Park:2012:EBB

[PTS^{+}12] Jongsoo Park, Ping Tak Peter Tang, Mikhail Smelyanskiy, Daehyun Kim, and Thomas Benson. Efficient backprojection-based synthetic aperture radar computation with many-core processors. In Hollingsworth [Hol12], pages 28:1–28:11. ISBN 1-4673-0804-

                                        8. URL http://conferences.
                                        computer.org/sc/2012/papers/
                                        1000a072.pdf.

Preissl:2012:CSS

[PWD^{+}12] Robert Preissl, Theodore M. Wong, Pallab Datta, Myron Flickner, Raghavendra Singh, Steven K. Esser, William P. Risk, Horst D. Simon, and Dharmendra S. Modha. Compass: a scalable simulator for an architecture for cognitive computing. In Hollingsworth [Hol12], pages 54:1–54:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a085.pdf.

Ravi:2012:VVB

[RBAC12] Vignesh T. Ravi, Michela Becchi, Gagan Agrawal, and Srimat Chakradhar. ValuePack: value-based scheduling framework for CPU–GPU clusters. In Hollingsworth [Hol12], pages 53:1–53:12. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a038.pdf.

Ravishankar:2012:CGP

[REP^{+}12] Mahesh Ravishankar, John Eisenlohr, Louis-Noël Pouchet, J. Ramanujam, Atanas Rountev, and P. Sadayappan. Code generation for parallel execution of a class of irregular loops on distributed memory systems. In Hollingsworth [Hol12], pages 72:1–72:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a100.pdf.

Riesen:2012:ASI

[RFD^{+}12] Rolf Riesen, Kurt Ferreira, Dilma Da Silva, Pierre Lemarinier, Dorian Arnold, and Patrick G. Bridges. Alleviating scalability issues of checkpointing protocols. In Hollingsworth [Hol12], pages 18:1–18:11. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a021.pdf.

Ren:2012:PWA

[RLY^{+}12] Yufei Ren, Tan Li, Dantong Yu, Shudong Jin, Thomas Robertazzi, Brian L. Tierney, and Eric Pouyoul. Protocols for wide-area data-intensive applications: design and performance issues. In Hollingsworth [Hol12], pages 34:1–34:11. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a059.pdf.

Rietmann:2012:FAS

[RMNM^{+}12] Max Rietmann, Peter Messmer, Tarje Nissen-Meyer, Daniel Peter, Piero Basini, Dimitri Komatitsch, Olaf Schenk, Jeroen Tromp, Lapo Boschi, and Domenico Giardini. Forward and adjoint simulations of seismic wave propagation on emerging large-scale GPU architectures. In Hollingsworth [Hol12], pages 38:1–38:11. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a104.pdf.

Sundar:2012:PGA

[SBB^{+}12] Hari Sundar, George Biros, Carsten Burstedde, Johann Rudi, Omar Ghattas, and Georg Stadler. Parallel geometricalgebraic multigrid on unstructured forests of octrees. In Hollingsworth [Hol12], pages 43:1–43:11. ISBN 1-4673-0804-

8. URL http://conferences.
computer.org/sc/2012/papers/
1000a045.pdf.

Schindewolf:2012:WSA

[SBG^{+}12] Martin Schindewolf, Barna Bihari, John Gyllenhaal, Martin Schulz, Amy Wang, and Wolfgang Karl. What scientific applications can benefit from hardware transactional memory? In Hollingsworth [Hol12], pages 90:1–90:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a073.pdf.

Satish:2012:LSE

[SKCD12] Nadathur Satish, Changkyu Kim, Jatin Chhugani, and Pradeep Dubey. Large-scale energy-efficient graph traversal: a path to efficient data-intensive supercomputing. In Hollingsworth [Hol12], pages 14:1–14:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a018.pdf.

Sharma:2012:DII

[SKYS12] Sushant Sharma, Dimitrios Katramatos, Dantong Yu, and Li Shi. Design and implementation of an intelligent end-to-end network QoS system. In Hollingsworth [Hol12], pages 68:1–68:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a077.pdf.

Sridharan:2012:SDF

[SL12] Vilas Sridharan and Dean Liberty. A study of DRAM failures in the field. In Hollingsworth [Hol12], pages 76:1–76:11. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a047.pdf.

Sarje:2012:MPX

[SLC^{+}12] Abhinav Sarje, Xiaoye S. Li, Slim Chourou, Elaine R. Chan, and Alexander Hexemer. Massively parallel X-ray scattering simulations. In Hollingsworth [Hol12], pages 46:1–46:11. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a067.pdf.

Sato:2012:DMN

[SMM^{+}12] Kento Sato, Naoya Maruyama, Kathryn Mohror, Adam Moody, Todd Gamblin, Bronis R. de Supinski, and Satoshi Matsuoka. Design and modeling of a non-blocking checkpointing system. In Hollingsworth [Hol12], pages 19:1–19:10. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a022.pdf.

Subramoni:2012:DSI

[SPK^{+}12] H. Subramoni, S. Potluri, K. Kandalla, B. Barth, J. Vienne, J. Keasler, K. Tomko, K. Schulz, A. Moody, and D. K. Panda. Design of a scalable InfiniBand topology service to enable networktopology-aware placement of processes. In Hollingsworth [Hol12], pages 70:1–70:12. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a076.pdf.

Speck:2012:MST

[SRK^{+}12] R. Speck, D. Ruprecht, R. Krause, M. Emmett, M. Minion, M. Winkel, and P. Gibbon. A massively space-time parallel Nbody solver. In Hollingsworth [Hol12], pages 92:1–92:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a083.pdf.

Spafford:2012:ADS

[SV12] Kyle L. Spafford and Jeffrey S.

Vetter. Aspen: a domain specific language for performance modeling. In Hollingsworth [Hol12], pages 84:1–84:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a091.pdf.

Sun:2012:OFG

[SZM^{+}12] Yanhua Sun, Gengbin Zheng, Chao Mei, Eric J. Bohm, James C. Phillips, Laximant V. Kalé, and Terry R. Jones. Optimizing fine-grained communication in a biomolecular simulation application on Cray XK6. In Hollingsworth [Hol12], pages 55:1–55:11. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a086.pdf.

Tang:2012:FLC

[TPKP12] Ping Tak Peter Tang, Jongsoo Park, Daehyun Kim, and Vladimir Petrov. A framework for low-communication 1-D FFT. In Hollingsworth [Hol12], pages 42:1–42:12. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a043.pdf.

Thorson:2012:SUF

[TW12] Greg Thorson and Michael Woodacre. SGI UV2: a fused computation and data analysis machine. In Hollingsworth [Hol12], pages 105:1–105:9. ISBN 1-4673-0804-8. URL

                                        http://conferences.computer.
                                        org/sc/2012/papers/1000a080.
                                        pdf.

VanderWijngaart:2012:EBN

[VSL12] Rob F. Van der Wijngaart, Srinivas Sridharan, and Victor W. Lee. Extending the BT NAS parallel benchmark to exascale computing. In Hollingsworth [Hol12], pages 94:1–94:9. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a094.pdf.

Wu:2012:RRA

[WHT^{+}12] Donghong Wu, Bingsheng He, Xueyan Tang, Jianliang Xu, and Minyi Guo. RAMZzz: rank-aware DRAM power management with dynamic migrations and demotions. In Hollingsworth [Hol12], pages 32:1–32:11. ISBN 1-4673-0804-

8. URL http://conferences.
computer.org/sc/2012/papers/
1000a042.pdf.

Williams:2012:OGM

[WKS^{+}12] Samuel Williams, Dhiraj D.

Kalamkar, Amik Singh, Anand M. Deshpande, Brian Van Straalen, Mikhail Smelyanskiy, Ann Almgren, Pradeep Dubey, John Shalf, and Leonid Oliker. Optimization of geometric multigrid for emerging multi- and manycore processors. In Hollingsworth [Hol12], pages 96:1–96:11. ISBN 1-4673-0804-

8. URL http://conferences.
computer.org/sc/2012/papers/
1000a095.pdf.

Wu:2012:HTM

[WLX^{+}12] Jingjin Wu, Zhiling Lan, Xuanxing Xiong, Nickolay Y. Gnedin, and Andrey V. Kravtsov. Hierarchical task mapping of cell-based AMR cosmology simulations. In Hollingsworth [Hol12], pages 75:1–75:10. ISBN 1-4673-0804-8. URL http:/

/conferences.computer.org/
sc/2012/papers/1000a035.pdf.

Xie:2012:COB

[XCD^{+}12] Bing Xie, Jeffrey Chase, David Dillow, Oleg Drokin, Scott Klasky, Sarp Oral, and Norbert Podhorszki. Characterizing output bottlenecks in a supercomputer. In Hollingsworth [Hol12], pages 8:1–8:11. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a007.pdf.

Zhang:2012:DAD

[ZKW^{+}12] Zhao Zhang, Daniel S. Katz, Justin M. Wozniak, Allan Espinosa, and Ian Foster. Design and analysis of data management in scalable parallel scripting. In Hollingsworth [Hol12], pages 85:1–85:11. ISBN 1-4673-0804-8. URL http:/

                                        /conferences.computer.org/
                                        sc/2012/papers/1000a016.pdf.